OpenStudy (anonymous):

Write the equation of the line with a slope of -1/3 and passing through the point (6, -4).

OpenStudy (goformit100):

Use 

OpenStudy (anonymous):

I would use point slope form because you have the slope and the points. y - y1 = m(x - x1) slope(m) = -1/3 (6,-4) x1 = 6, y1 = -4 now sub these into your formula

OpenStudy (anonymous):

I know how to do all of this.. but the answer that I'm getting is apparently wrong?

OpenStudy (anonymous):

y - y1 = m(x - x1) y - (-4) = -1/3(x - 6) y + 4 = -1/3x + 2 y = -1/3x + 2 - 4 y = -1/3x - 2

OpenStudy (anonymous):

oops...sorry, I lost track of time....I got to go now

OpenStudy (anonymous):

I ended up getting that answer. Thank you!

OpenStudy (anonymous):

your welcome :)